Suppose cost of production of a firm is given by:


TC=128+50Q-16Q^2+2Q^3


What amount of the product should be produced to maximize profit of the firm if price per unit output is 74? Find the maximum profit or loss.



1
Expert's answer
2022-02-13T11:50:24-0500

Profit is maximized, when P = MC.

"MC = TC'(Q) = 50 - 32Q + 6Q^2 = 74."

"3Q^2 - 16Q - 12 = 0,"

Q = (16 + 20)/6 = 6 units.

"TP = TR - TC = 74\u00d76 - (128+50\u00d76-16\u00d76^2+2\u00d76^3) = 160."